# AniFaceDiff: High-Fidelity Face Reenactment via Facial Parametric Conditioned Diffusion Models

### v1

@@ 3.3 Pose and Expression Conditioning
Our pose and expression conditioning module extracts pose and expression from the driving video, and identity-related information from the source image based on a pretrained off-the-shelf 3D face model, DECA. Subsequently, we condition the diffusion model using two components: improved spatial conditioning and non-spatial conditioning (the expression adapter).
Improved Spatial Conditioning with Facial Shape Alignment
@@
We utilize the encoder of DECA to estimate facial parameters from each frame of the driving video directly. Subsequently, we render these parameters into a sequence of 2D facial surface normal snapshots. However, in addition to containing the pose and expression information we require, these snapshots also include identity information reflected by the facial geometry from the driving frames. The identity information essential for our purpose should be extracted from the source image. Any inconsistency in the facial geometry information between the condition and the final output can result in sub-optimal performance, particularly noticeable when there is a significant mismatch in identity between the source image and the driving video. To address this issue, we utilize a facial shape alignment strategy which combines the identity information from the source image with the pose and expression information from the driving frames to generate the refined 2D facial conditions.
@@
Specifically, we obtain shape parameters $\beta$ containing identity information from the source image, and other parameters including pose $\theta$ and expression $\psi$ from the driving frames by the DECA encoder. We then use FLAME to generate the 3D face meshes and render them into surface normal snapshots to eliminate the effects of lighting conditions. The entire process can be represented as:
$n^{1:N}=\mathcal{R}(M(\beta_{s},\theta^{1:N}_{d},\psi^{1:N}_{d}),c^{1:N}_{d}),$ (2)
@@
where $n^{1:N}$ represents the 2D facial conditions, the rendered surface normal snapshots, for frames $1:N$. $\mathcal{R}$ represents the rasterizer (PyTorch3D [60]). M represents FLAME model used to obtain head vertices with facial parameters. $c$ represents the camera information to project the 3D mesh into image space. $\beta_{s}$ is extracted from the source image, while $\theta^{1:N}_{d}$, $\psi^{1:N}_{d}$ and $c^{1:N}_{d}$ are extracted from the driving frames.
@@
After extracting the 2D facial conditions which are spatial-aligned with the output frames, we encode them with the Pose Encoder, a series of convolution layers, similar to Animate Anyone [53], to project the conditions to the same resolution with the noisy input of denoising UNet. Subsequently, we concatenate the spatial-aligned condition with the noisy input and feed it to the denoising UNet.
@@
Crucially, the low resolution of the 3D face mesh leads to the loss of mid-frequency details related to expression. Drawing inspiration from IP-Adapter [46] for image prompt conditioning, we introduce an expression adapter conditioning mechanism to address potential information loss resulting from spatial modulation, as illustrated in the lower branch of Fig. 1. Unlike IP-Adapter for conditional image generation, we integrate video-level expression prompts from the driving frames in parallel with the image prompt from the source image. Specifically, we extract the expression embedding $\mathbf{\psi}\in\mathbb{R}^{N\times 50}$ from N driving frames using DECA and project it into the same dimension as the CLIP image embedding. This projected expression embedding $\mathit{p_{\psi}}$ is then injected into the diffusion model through additional cross-attention layers, which can be represented as:
$Z_{new}=Attention(Q,K^{i},V^{i})+\lambda\cdot Attention(Q,K^{\psi},V^{\psi}),$ (3)
where $\mathit{Q}$, $\mathit{K^{i}}$, and $\mathit{V^{i}}$ are the query, key, and value matrices from the original cross-attention layer for CLIP image features, while $\mathit{K^{\psi}}$ and $\mathit{V^{\psi}}$ are from the additional cross-attention layer for expression features. The projected expression embedding $\mathit{p_{\psi}}$ is injected into the cross-attention layer as $\mathit{K^{\psi}=p_{\psi}W_{k}^{\psi}}$ and $\mathit{V^{\psi}=p_{\psi}W_{v}^{\psi}}$, where $\mathit{W_{k}^{\psi}}$ and $\mathit{W_{v}^{\psi}}$ are the corresponding weight matrices. In the 3D denoising UNet for generating a video, the CLIP embedding of the source image is repeated N times to match the length of the expression embedding from the driving frames.
@@ 3.4 Training and Inference
The training process comprises two stages. In the first stage, the model is trained to generate a single image under the guidance of a source image and a driving frame. During the training process, the model utilizes same-identity construction as its objective, where both the source image and the driving frame are derived from the same video. DECA is utilized to extract facial parameters and render facial snapshots from the driving frame. The driving frame $\textbf{x}_{0}$ is first encoded into the latent representation $\textbf{z}_{0}$, which is then noised with $\epsilon\sim\mathcal{N}(0,1)$ at timestep t via a defined scheduler as:
@@
where $\bar{\alpha}_{t}={\textstyle\prod_{s=0}^{t}}\alpha_{s}$, $\alpha_{t}$ is the variance schedule, and $\textbf{z}_{t}$ is the noisy latent. The model is trained to predict the added noise by denoising function $\epsilon_{\theta}$ with condition features f from the source image and the driving video. The VAE, CLIP image encoder, and DECA are pretrained and kept frozen while the pose encoder, the expression adapter, the ReferenceNet, and the denoising UNet are updated. The training objective of the first stage is similar to that of Stable Diffusion [1]:
@@
In the second stage, the models are trained to generate a temporally consistent video clip (N frames) with guidance from both a source image and a driving video clip. Pretrained temporal modules [59] are integrated into the denoising UNet. During this stage, only the temporal modules are updated and all other components remain frozen. The training objective of the second stage is formulated as:
@@
At inference, we utilize DECA to extract information from both the source image and the driving video. Leveraging a facial shape alignment strategy, we refine the facial snapshot sequence. Begining with noises $\textbf{z}_{t}^{1:N}$ sampled from a Gaussian distribution in the latent space, the model conduct denoising process to $\textbf{z}_{t-1}^{1:N}$ by predicting noise using the denoising UNet with the condition $\textbf{f}^{1:N}$ and iteratively to $\textbf{z}_{0}^{1:N}$ via a defined sampling process. Finally, the denoised outputs $\textbf{z}_{0}^{1:N}$ are projected back to the pixel level $\textbf{x}_{0}^{1:N}$ to obtain the generated video via the VAE decoder.
@@ 4 Experiments
In our experiments, we utilized the VoxCeleb dataset [2] after preprocessing as described in [5]. This dataset comprises 19,522 and 525 utterance sequences, spanning 453 and 13 identities, with 3,354 and 90 videos in the training and test sets, respectively.
@@
Training was conducted on $256\times 256$ images using 2 NVIDIA A100 GPUs. In the first stage, we trained for over 100,000 steps with a batch size of 32. Subsequently, in the second stage, we trained for over 20,000 steps using a batch size of 8 for a video clip consisting of 24 frames. All models were optimized using the Adam optimizer with a learning rate of $1\times 10^{-5}$.
4.1 Comparisons
@@
We compare our method with four state-of-the-art methods, namely FOMM [5], Face vid2vid [6], TPSMM [7], and HyperReenact [15]. Results for the baseline methods are generated using their provided code and pretrained weights, with the exception of Face vid2vid, for which we utilize unofficial code [^1] https://github.com/zhanglonghao1992/One-Shot_Free-View_Neural_Talking_Head_Synthesis due to the absence of the official version. We assess our method and the baseline approaches using four distinct metrics. First, we utilize the Fréchet-Inception Distance (FID) to evaluate image quality by measuring the distribution discrepancy between the generated data and the test set of VoxCeleb. Second, we employ the identity preservation cosine similarity (CSIM), following the approach outlined in Encoding in Style [61], to quantify the preservation of identity. Third, we measure the accuracy of pose and expression transfer using the Average Pose Distance (APD) and the Average Expression Distance (AED). CSIM is computed based on the source image and the outputs, while APD and AED are calculated based on the driving video and the outputs.
Quantitative Comparisons
@@
. Same-identity Reconstruction Cross-identity Reenactment FID CSIM APD AED FID CSIM APD AED User Pref. (%) FOMM [5] 21.91 0.83 0.011 0.092 41.93 0.51 0.028 0.210 10.25 Face vide2vid [6] 18.57 0.84 0.012 0.092 35.21 0.56 0.034 0.238 9.50 TPSMM [7] 20.18 0.84 0.009 0.081 38.91 0.52 0.024 0.191 16.50 HyperReenact [15] 72.72 0.48 0.011 0.098 73.35 0.40 0.017 0.181 20.75 Ours 17.21 0.75 0.011 0.091 32.58 0.57 0.022 0.174 43.00
@@
We evaluate our method and the baselines on the test set of VoxCeleb. In the evaluation, higher scores indicate better performance for CSIM, while lower scores are preferable for all other metrics. We conduct one-shot face reenactment on two tasks: same-identity reconstruction, which mirrors the training process, and the more challenging primary task of cross-identity reenactment.
@@
For the same-identity reconstruction task, we randomly selected 100 videos from the VoxCeleb test set, using the first frame as the source image and considering the remaining frames as the driving frames. As shown in Table 1, our method achieves the best image quality (FID) and ranks second for pose (APD) and expression accuracy (AED). TPSMM outperforms all other methods in terms of CSIM, APD, and AED. All warping-based methods (FOMM, Face vid2vid, and TPSMM) perform well on this proxy task when there is no identity variance and usually a small pose difference between the source image and the driving video.
@@
Even though current methods perform well on same-identity reconstruction, the main task (cross-identity reenactment), where the source image and driving video are from different identities, introduces substantial identity mismatches and pose variations, demanding higher generalization capabilities from models. Similar to previous work, we randomly select 100 image-video pairs from the VoxCeleb test set. Table.1 illustrates a significant performance drop for each method when transitioning from same-identity reconstruction to cross-identity reenactment. As observed in previous SOTA methods, there is a trade-off between identity consistency (CSIM) and expression accuracy (AED). The more modifications related to expression made to the original source identity, the less the generated image will resemble the original, as assessed by face recognition models. Despite this trade-off, Our method (AniFaceDiff) achieves SOTA results on both CSIM and AED. Additionally, we outperform all four baseline methods in terms of image quality (FID) and secure the second position in terms of pose accuracy (APD).
Qualitative Comparisons
Figure 2: Qualitative comparison with SOTA methods on cross-identity reenactment. Our method achieves high identity preservation and expression accuracy. (All photorealistic human images presented in this paper are virtual individuals generated by [33] and do not exist in the real world.)
Figure 3: Qualitative comparison with SOTA methods on cross-identity reenactment in the presence of significant pose variations between the source and driving images. Our method accurately captures the source identity and faithfully transfers the target pose and expression, even under significant pose variations.
We present a few representative examples of cross-identity reenactment in Fig. 2. Additional examples can be found in Appendix A.2. Overall, our method generates high quality images compared to all other state-of-the-art methods while preserving both source identity and pose and expression of the driving frame. Better preservation of source identity is especially evident in the first and second rows, where other methods show noticeable influence from the driving identity. In contrast, our method maintains the source identity with greater accuracy. Furthermore, our approach excels at maintaining the expressions and poses of the driving videos, as seen in the third and fourth rows, where our method consistently captures expression-related details such as wrinkles, brow furrows and lip curvature more accurately than other state-of-the-art methods.
Fig.3 demonstrates a more challenging scenario where there are significant pose differences between the source and driving images. FOMM, Face vid2vid, and TPSMM tend to produce blurry outputs with severe artifacts and distortions due to their reliance on the optical flow field. Additionally, their inadequate inpainting capabilities result in blurry and unrealistic regions when generating areas not present in the source image. HyperReenact, which does not rely on the optical flow field, generates more fine-grained results. However, it still suffers from a loss of identity and detailed information, such as facial wrinkles, hair texture, and lighting. This is likely due to the limitations of GAN inversion and the use of a pretrained ArcFace [62] for extracting identity information from the source image. In contrast, our method excels in preserving detailed information from the source image, while accurately maintaining pose and expression from the driving video due to the designed conditioning module. Overall, our method produces the cleanest and most high-fidelity results among all the state-of-the-art methods, even when the source face is partially occluded (row 3) or when there is a large distance difference from the camera between the source and driving images (row 4).
We note that the current metrics might not fully capture the qualitative differences observed in the results. Hence, we conducted a user study involving 20 participants to further assess performance based on human visual experience. We presented 20 randomly selected source-driving image pairs: 5 for same-identity reconstruction and 15 for cross-identity reenactment, as cross-identity reenactment is the primary task. We provided users with three evaluation criteria similar to those used by HyperReenact [15]: 1) image quality, 2) identity and nuanced detail preservation from the source image, and 3) pose and expression preservation from the driving image. As shown in Table 1, our method significantly outperforms all other SOTA methods based on user preference.
Out-of-domain Data
Figure 4: Qualitative comparison with SOTA methods on out-of-domain (non-photorealistic) face reenactment. Our method can generate high-quality images even without being trained on such types of data. (All out-of-domain facial images are generated by [63])
We also compare our method with these SOTA methods outside the training distribution to evaluate generalization capabilities. As shown in Fig. 4, The first and second rows demonstrate out-of-domain (non-photorealistic) source images reenacted by out-of-domain driving images. The third and fourth rows demonstrate out-of-domain source images reenacted by in-domain (photorealistic) driving images. Our method exhibits strong generalization capabilities on out-of-domain data. More results can be found in Appendix A.3.

### v2

@@ 3.3 Pose and Expression Conditioning
Our pose and expression conditioning module extracts pose and expression from the target video, and identity-related information from the reference image based on a pretrained off-the-shelf 3D morphable model, DECA. Subsequently, we condition the diffusion model using two components: improved spatial conditioning and non-spatial conditioning (the expression adapter).
Improved Spatial Conditioning with Facial Alignment
@@
We utilize the encoder of DECA to estimate head parameters from each frame of the target video directly. Subsequently, we render these parameters into a sequence of 2D surface normal snapshots for spatial-aligned conditioning. However, in addition to containing the pose and expression information we require, such snapshots inherently contain the identity characteristic reflected by the head geometry of the motion frame rather than that of the reference image. The inconsistency in the facial geometry between the spatial-aligned condition and the final output can result in sub-optimal performance, especially when there is a significant mismatch in identity characteristics between the reference image and the target video—a challenge that is particularly pronounced in stylized animation. Additionally, when there is a substantial head scale (the relative size or proportions of the face within the image) discrepancy between the reference image and the motion frames, the spatial information provided by the snapshots may conflict with that provided by the ReferenceNet. To address these issues, we utilize a strategy named Facial Alignment, which consists of two components: shape alignment, which integrates head shape from the reference image with pose and expression information from the target frames, and scale alignment, which ensures alignment of the 2D surface normal snapshots with the scale of the reference head.
@@
Specifically, we obtain shape parameters $\beta$ containing identity characteristic from the reference image, and other parameters including pose $\theta$ and expression $\psi$ from the target frames by the DECA encoder. The decoding process consists of generating 3D head vertices based on FLAME, applying the similarity transformation to head vertices using the rescaled transformation matrices, and render them into surface normal snapshots to eliminate the effects of lighting conditions. The entire process can be represented as:
$n^{1:N}=\mathcal{R}(\mathcal{T}(\mathcal{F}(\beta_{s},\theta^{1:N}_{d},\psi^{1:N}_{d}),c^{1:N}_{d}),a^{1:N}),$ (2)
@@
where $n^{1:N}$ represents the 2D head conditions, the rendered surface normal snapshots, for frames $1:N$. $\mathcal{R}$ represents the rasterizer (PyTorch3D [58]). $\mathcal{F}$ represents FLAME model used to obtain head vertices with facial parameters. $c$ represents the camera information to project the 3D mesh into image space. $\beta_{s}$ is extracted from the reference input, while $\theta^{1:N}_{d}$, $\psi^{1:N}_{d}$ and $c^{1:N}_{d}$ are extracted from the target motion frames. $\mathcal{T}$ represents the similarity transformation with the rescaled matrices $a^{1:N}$. We derive a rescaled similarity transformation matrix for each frame, using a rescaling factor computed from the facial bounding boxes of the reference image and the first motion frame
@@
After extracting the 2D surface normal snapshots which are spatial-aligned with the output frames, we encode them with the Pose Encoder, a series of convolution layers, similar to Animate Anyone [50], to project the conditions to the same resolution with the noisy input of denoising UNet. Subsequently, we add the spatial-aligned condition to the noisy input and feed it to the denoising UNet.
@@
Crucially, the low resolution of 3D face meshes results in the loss of mid-frequency expression details, which are essential for producing high-quality expressive stylized animations. Drawing inspiration from IP-Adapter [43] for image prompt conditioning, we introduce an expression adapter conditioning mechanism to address potential information loss resulting from spatial modulation, as illustrated in the lower branch of Fig. 2. Unlike IP-Adapter, which is used for conditional image generation, we incorporate video-level expression prompts (DECA expression embeddings) from the target motion frames in parallel with the image prompt (CLIP embedding) from the reference image. This module is designed to enhance the transfer of expressions while preserving the semantic information from the reference image. Specifically, we extract the expression embedding $\mathbf{\psi}^{1:N}\in\mathbb{R}^{50}$ from N motion frames using DECA and project it into the same dimension as the CLIP reference embedding. This projected expression embedding $\mathit{p_{\psi}^{1:N}}$ is then injected into the diffusion model through additional cross-attention layers, which can be represented as:
$Z_{new}=Attention(Q,K_{i},V_{i})+\lambda\cdot Attention(Q,K_{\psi},V_{\psi}),$ (3)
where $\mathit{Q}$, $\mathit{K_{i}}$, and $\mathit{V_{i}}$ are the query, key, and value matrices from the cross-attention layers for CLIP image features, while $\mathit{K_{\psi}}$ and $\mathit{V_{\psi}}$ are from the additional cross-attention layer for expression features. The projected expression embedding $\mathit{p_{\psi}^{1:N}}$ is injected into the cross-attention layer as $\mathit{K_{\psi}=p_{\psi}^{1:N}W^{k}_{\psi}}$ and $\mathit{V_{\psi}=p_{\psi}^{1:N}W^{v}_{\psi}}$, where $\mathit{W_{k}^{\psi}}$ and $\mathit{W_{v}^{\psi}}$ are the corresponding weight matrices. In the 3D denoising UNet for generating a video, the CLIP embedding of the reference image is repeated N times to match the length of the expression embedding from the motion frames.
@@ 3.4 Training and Inference
The training process comprises two stages. In the first stage, the model is trained to generate a single image under the guidance of a reference image and a target motion frame. During the training process, the model utilizes same-identity construction as its objective, where both the reference image and the motion frame are derived from the same video. DECA is utilized to extract head parameters and render snapshots from the motion frame. The target motion frame $\textbf{x}_{0}$ is first encoded into the latent representation $\textbf{z}_{0}$, which is then noised with $\epsilon\sim\mathcal{N}(0,1)$ at timestep t via a defined scheduler as:
@@
where $\bar{\alpha}_{t}={\textstyle\prod_{s=0}^{t}}\alpha_{s}$, $\alpha_{t}$ is the variance schedule, and $\textbf{z}_{t}$ is the noisy latent. The model is trained to predict the added noise by denoising function $\epsilon_{\theta}$ with condition features f from the reference image and the target video. The VAE, CLIP image encoder, and DECA are pretrained and kept frozen while the pose encoder, the expression adapter, the ReferenceNet, and the denoising UNet are updated. The training objective of the first stage is similar to that of Stable Diffusion [1]:
@@
In the second stage, the models are trained to generate a temporally consistent video clip (N frames) with guidance from both a reference image and a target motion video clip. Pretrained temporal modules [57] are integrated into the denoising UNet. During this stage, only the temporal modules are updated and all other components remain frozen. The training objective of the second stage is formulated as:
@@
At inference, we utilize DECA to extract information from both the reference image and the target video. Leveraging the facial alignment strategy, we refine the surface normal snapshot sequence. Beginning with noises $\textbf{z}_{t}^{1:N}$ sampled from a Gaussian distribution in the latent space, the model conduct denoising process to $\textbf{z}_{t-1}^{1:N}$ by predicting noise using the denoising UNet with the condition $\textbf{f}^{1:N}$ and iteratively to $\textbf{z}_{0}^{1:N}$ via a defined sampling process. Finally, the denoised outputs $\textbf{z}_{0}^{1:N}$ are projected back to the pixel level $\textbf{x}_{0}^{1:N}$ to obtain the generated video via the VAE decoder.
@@ 4 Experiments
4.1 Data and Training Details
In our experiments, we generated a diverse array of styles (e.g. artistic, cartoon, sculpture, and painting) using [2] to conduct out-of-domain animation experiments. This broad spectrum of styles enabled us to evaluate the robustness and adaptability of our method across a variety of domains beyond the original dataset distribution. We utilized the subset of VoxCeleb dataset [59] after preprocessing as described in [5] for training and test. We also test the methods on the test set of VoxCeleb2 dataset [60].
@@ 4.1 Data and Training Details
Training was conducted on $256\times 256$ images using 2 NVIDIA A100 80G GPUs for around 4 days. In the first stage, we trained for over 100,000 steps with a batch size of 32. Subsequently, in the second stage, we trained for over 20,000 steps using a batch size of 8 for a video clip consisting of 24 frames. All models were optimized using the Adam optimizer with a learning rate of $1\times 10^{-5}$. During the inference phase, our method generates results based on a given reference image and a target motion video, utilizing the DDIM sampler [61] with 20 denoising steps. The entire process takes approximately 25 seconds to produce a 100-frame video clip using a single NVIDIA A100 80G GPU.
4.2 Out-of-Domain Animation
Figure 3: Qualitative comparison with SOTA methods on out-of-domain stylized animation. Our method can generate high-quality images even without explicit training on such types of data.
Figure 4: Qualitative comparison with SOTA methods on out-of-domain stylized animation in the presence of significant pose variations between reference images and target motion. Our method accurately captures the reference features and faithfully realize the target pose and expression, even under significant pose variations.
We compare our method with four state-of-the-art methods, namely FOMM [5], Face vid2vid [6], TPSMM [7], and HyperReenact [26]. Results for the baseline methods are generated using their provided code and pretrained weights, with the exception of Face vid2vid, for which we utilize unofficial code [^1] https://github.com/zhanglonghao1992/One-Shot_Free-View_Neural_Talking_Head_Synthesis due to the absence of the official version. We compare our method with state-of-the-art (SOTA) methods on out-of-domain data to assess its generalization capabilities across a diverse range of styles. As shown in Fig. 3, our method demonstrates strong generalization on out-of-domain data, successfully animating characters from various styles and domains. Notably, stylized avatars can be animated using human-like virtual faces (as shown in row 1 and row 2 of both Fig. 3 and Fig. 4) as well as stylized avatars from other domains. Our method preserves the appearance and intricate details of the reference image across different styles while accurately capturing and generating the pose and expression from the motion video even without being trained on such data. In contrast, warping-based methods (FOMM, Face vid2vid, and TPSMM) exhibit noticeable identity leakage when there is a significant head shape difference between the reference and target images. HyperReenact, on the other hand, fails to preserve both the style and appearance of the reference image. This is likely due to the limitations of GAN inversion and the use of a pre-trained ArcFace model [62]. This strong generalization capability makes our method highly adaptable for various applications, including animation production, privacy protection, virtual education, and beyond. It enables the animation of virtual avatars in diverse styles, offering flexibility and enhanced user experience. Furthermore, the ability to handle out-of-domain data addresses privacy and ethical concerns, as stylized or non-photorealistic avatars reduce the potential risks of misinterpretation or misuse.
Fig. 4 illustrates a more challenging scenario involving significant pose differences between the reference and target images. In this case, FOMM, Face vid2vid, and TPSMM tend to produce blurry outputs with severe artifacts and distortions due to their reliance on optical flow fields. The blurriness is particularly noticeable in regions where motion changes occur. Moreover, these methods struggle with inpainting, resulting in blurry and unnatural regions when generating areas not present in the reference image. HyperReenact, which does not depend on optical flow fields, generates finer-grained results compared to the warping-based methods. However, it still suffers from a loss of style, appearance, and fine details, such as hair texture and lighting. In contrast, our method excels in preserving detailed information from the reference input while accurately maintaining pose and expression from the target motion video, thanks to the specially designed conditioning module. Expression-related details, such as wrinkles, brow furrows, and lip curvature, can be faithfully generated. Overall, our method produces the cleanest, most high-fidelity results among all the state-of-the-art methods, even when only the side profile is available (as shown in row 1) or the head is partially occluded (as shown in row 4).
4.3 Within-Domain Animation
@@
We assess our method and the baseline approaches using five distinct metrics on within-domain animation. First, we utilize the Fréchet-Inception Distance (FID) to evaluate image quality by measuring the distribution discrepancy between the generated data and the test set and the Fréchet Video Distance (FVD) to evaluate video quality. Both FID and FVD implementations are based on StyleGAN-V [63]. Second, we employ the identity preservation cosine similarity (CSIM), following the approach outlined in Encoding in Style [64]. Third, we measure the accuracy of pose and expression using the Average Pose Distance (APD) and the Average Expression Distance (AED). CSIM is computed based on the reference image and the outputs, while APD and AED are calculated based on the target video and the outputs.
@@
. Same-identity Reconstruction Cross-identity Animation FID CSIM APD AED FVD FID CSIM APD AED FVD FOMM [5] 26.48 0.83 0.011 0.092 142.18 50.36 0.51 0.028 0.210 254.60 Face vide2vid [6] 23.43 0.84 0.012 0.092 136.46 43.23 0.56 0.034 0.238 210.34 TPSMM [7] 25.44 0.84 0.009 0.081 146.98 44.66 0.52 0.024 0.191 253.44 HyperReenact [26] 98.33 0.48 0.011 0.098 379.07 98.92 0.40 0.017 0.181 364.83 Ours 21.25 0.75 0.011 0.091 168.76 36.26 0.57 0.022 0.174 241.83
Table 2: Quantitative results on VoxCeleb2. The best results are shown in bold, and the second-best results are shown in underline
. Same-identity Reconstruction Cross-identity Animation FID CSIM APD AED FVD FID CSIM APD AED FVD FOMM [5] 43.53 0.77 0.019 0.124 178.42 59.67 0.47 0.035 0.235 242.93 Face vide2vid [6] 40.91 0.78 0.021 0.124 181.64 51.13 0.54 0.039 0.232 238.19 TPSMM [7] 45.52 0.76 0.018 0.109 196.89 53.39 0.48 0.032 0.231 251.45 HyperReenact [26] 101.07 0.48 0.013 0.101 380.31 97.72 0.42 0.016 0.163 405.88 Ours 34.46 0.67 0.016 0.119 185.54 33.81 0.55 0.024 0.188 233.91
@@
We evaluate our method and the baselines on the test set of VoxCeleb. In the evaluation, higher scores indicate better performance for CSIM, while lower scores are preferable for all other metrics. We conduct one-shot animation on two tasks: same-identity reconstruction, which mirrors the training process, and the more challenging primary task of cross-identity animation.
@@
For the same-identity reconstruction task, we randomly selected 100 videos from the VoxCeleb test set, using the first frame as the reference image and considering the remaining frames as the motion frames. As shown in Table 1, our method achieves the best image quality (FID) and ranks second for pose (APD) and expression accuracy (AED). TPSMM outperforms all other methods in terms of CSIM, APD, and AED. All warping-based methods (FOMM, Face vid2vid, and TPSMM) perform well on this proxy task when there is no appearance variance and usually a small pose difference between the reference image and the target motion.
@@
Even though current methods perform well on same-identity reconstruction, the main task (cross-identity animation) introduces substantial appearance mismatches and pose variations, demanding higher generalization capabilities from models. Similar to previous work, we randomly select 100 image-video pairs from the VoxCeleb test set. Table.1 illustrates a significant performance drop for each method when transitioning from same-identity reconstruction to cross-identity animation. As observed in previous SOTA methods, there is a trade-off between identity consistency (CSIM) and expression accuracy (AED). The more modifications related to expression made to the original reference image, the less the generated identity characteristics will resemble the original, as assessed by facial models. Despite this trade-off, Our method (AniFaceDiff) achieves SOTA results on both CSIM and AED. Additionally, our method outperforms all four baseline methods in terms of image quality (FID) and secures the second position in terms of pose accuracy (APD) and video quality (FVD).
To assess the generalization ability, we tested both our method and other state-of-the-art approaches on the VoxCeleb2 test set (all models were trained on VoxCeleb). Similar to the evaluation protocol used for VoxCeleb, for same-identity reconstruction, we randomly selected 100 videos from the VoxCeleb2 test set, using the first frame as the reference image and the remaining frames as motion frames. For cross-identity animation, we randomly selected 100 pairs from the test set. The results are presented in Table 2. Consistent with the VoxCeleb evaluation, our method dominates in FID for same-identity reconstruction. For cross-identity animation, it significantly outperforms other approaches in FID and achieves the top scores in both CSIM and FVD, demonstrating strong generalization ability. Additionally, our method ranks second in APD and AED.